OpenStudy (anonymous):

"Determine the best method to solve the system of equations. Then solve the system. 5x-2y=7 2x+5y=-3" Solve please?

OpenStudy (mertsj):

Multiply top equation by 2 and bottom equation by negative 5

OpenStudy (mertsj):

10x-4y=14 -10x-25y=15 -29y=29 y=-1 Plug that in and find x

OpenStudy (anonymous):

Thanks!